Show your solution and provide units in your solution

1.Calculate the mass(in grams) of magnesium chloride (MgCl2) that would be needed to prepare 150 mL of a 20% by mass aqueous solution of the salt.The density of the solution is 1.1g/mL.

3. A solution is prepared by dissolving 20 mL of pure hydrogen peroxide (H2O2) in enough water to make 200 mL of solution.What is the concentration of the H2O2 solution?


1
Expert's answer
2021-10-15T02:55:30-0400
  1. Mass of solution=volume*Density

=1.1g/ml(150ml)=165g

Mass of MgCl2=(20/100)*165g

=33g of MgCl2

2.Concentration of pure hydrogen peroxide=42.404g/L

Molarity=(g/l)/(RFM)

=(42.404g/L)/34g

=1.2472M

M1V1=M2V2

20*1.2472=200M2

M2=(20*1.2472)/200

=0.1247M H2O2